# Enhancement of superconductivity on thin film of Sn under high pressure

### v1

@@ Abstract
We investigated the pressure effects of a superconductivity on thin films of Sn. Elemental superconductor Sn with a body-centered tetragonal structure, -Sn, exhibits superconductivity below the superconducting transition temperature ( K) at ambient pressure. of Sn increases with lowering dimension such as in thin film and nanowire growth, or by high-pressure application. For thin films, exhibits a slight increase up to approximately 4 K compared to the bulk value, attributable to the crystalline size and lattice disorder. By applying pressure on a bulk Sn, decreases with increasing pressure from 3.72 K to 5.3 K with the structural transformation into -Sn around 10 GPa. However, the combination of these effects on thin films of Sn, namely, thin-film growth and pressure effects, remains underexplored. In this study, we combined film-growth and pressure-application techniques to further increase using a diamond anvil cell with boron-doped diamond electrodes. The drop of the electrical resistance suggesting the onset of on the thin film reached above 6 K in -Sn phase. Further, the upper critical magnetic field was drastically enhanced. Atomic force microscopy suggests that the refinement of the grain size of the thin film under the non-hydrostatic pressure conditions contributes to stabilizing the higher of -Sn.
@@ III.3 Non-hydrostatic pressure on thin film
Figures 4(a) and (b) show the dependence of for the thin film of Sn under various at 2.5 and 10.5 GPa under non-hydrostatic pressure condition. was observed at 1 T, whereas the upper critical magnetic field () was approximately 0.1 T under ambient pressure. Drastic enhancement was also observed in the phase at 10.5 GPa. As shown in Fig. 4(c), estimated using Werthamer–Helfand–Hohenberg (WHH) model reached several teslas at high pressures Helfand and Werthamer (1966); Werthamer et al. (1966); Baumgartner et al. (2013). Notably, the enhancement of -Sn under the non-hydrostatic pressure condition was observed reproducibly. Figure 5(b) shows the dependence of for the thin film of Sn with the other setup at 5 GPa. The optical image of the thin film of Sn is shown in Fig. 5(a). The decrease suggesting the superconducting transition was observed even at T. The enhancement of is thought to originate from the shortening in the mean free path of the electrons and the resulting shortening in the coherence length. The RRR was estimated to be 11 under ambient pressure; however, it decreased to 1.4 under non-hydrostatic pressure condition, indicating that the mean free path of the electrons shortened.
@@ III.4 Quasi-hydrostatic pressure on thin film
We performed high-pressure generation with better hydrostatic pressure condition using a liquid pressure transmitted medium, glycerol. Figure 6(a) shows the – of the thin film at 9 GPa under quasi-hydrostatic pressure conditions. The optical image of the thin film inside the DAC is shown in the inset of Fig. 6(a). The RRR was with in 2–-3 in quasi-hydrostatic pressure measurements, which was slightly higher than those for the non-hydrostatic pressure condition. The drop suggesting superconducting transition was observed around 5 K. Figure 6(b) shows – under quasi-hydrostatic pressure condition between 2 and 8 K with warming process. The value was normalized using the value at 8 K. At 9 GPa, slightly decreased with decreasing around 5.3 K, suggesting the superconducting transition on -Sn. The drop was also observed around 3 K. The -Sn phase became significant on further pressure application. The onset of slightly increased at 12 GPa. We also decreased the pressure from 12 to 8 GPa. The decrease suggesting the superconducting transition of -Sn was slightly observed, whereas the -Sn was clearly observed at 8.5 GPa under the non-hydrostatic pressure condition. With further pressure decrease, -Sn vanished at 2 GPa.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant No. 19H02177, 20H05644 and JST-Mirai Program Grant No. JPMJMI17A2. This work was also supported by the Joint Usage/Research Center PRIUS, Ehime University, JapanReferences
- Mao et al. (2018) H.-K. Mao, X.-J. Chen, Y. Ding, B. Li, and L. Wang, Rev. Mod. Phys. 90, 015007 (2018).
- Yamanaka (2010) S. Yamanaka, Dalton Trans. 39, 1901 (2010).
- Shimizu et al. (1998) K. Shimizu, K. Suhara, M. Ikumo, M. I. Eremets, and K. Amaya, Nature 393, 767 (1998).
- Shimizu et al. (2005) K. Shimizu, K. Amaya, and N. Suzuki, J. Phys. Soc. Jpn. 74, 1345 (2005).
- Drozdov et al. (2015) A. P. Drozdov, M. I. Eremets, I. A. Troyan, V. Ksenofontov, and S. I. Shylin, Nature 525, 73 (2015).
- Drozdov et al. (2019) A. P. Drozdov, P. P. Kong, V. S. Minkov, S. P. Besedin, M. A. Kuzovnikov, S. Mozaffari, L. Balicas, F. F. Balakirev, D. E. Graf, V. B. Prakapenka, E. Greenberg, D. A. Knyazev, M. Tkacz, and M. I. Eremets, Nature 569, 528 (2019).
- Somayazulu et al. (2019) M. Somayazulu, M. Ahart, A. K. Mishra, Z. M. Geballe, M. Baldini, Y. Meng, V. V. Struzhkin, and R. J. Hemley, Phys. Rev. Lett. 122, 027001 (2019).
- Matsumoto et al. (2020) R. Matsumoto, M. Einaga, S. Adachi, S. Yamamoto, T. Irifune, K. Terashima, H. Takeya, Y. Nakamoto, K. Shimizu, and Y. Takano, Supercond. Sci. Technol. 33, 124005 (2020).
- Semenok et al. (2020) D. V. Semenok, A. G. Kvashnin, A. G. Ivanova, V. Svitlyk, V. Y. Fominski, A. V. Sadakov, O. A. Sobolevskiy, V. M. Pudalov, I. A. Troyan, and A. R. Oganov, Materials Today 33, 36 (2020).
- Chen et al. (2021) W. Chen, D. V. Semenok, X. Huang, H. Shu, X. Li, D. Duan, T. Cui, and A. R. Oganov, Phys. Rev. Lett. 127, 117001 (2021).
- Kong et al. (2021) P. Kong, V. S. Minkov, M. A. Kuzovnikov, A. P. Drozdov, S. P. Besedin, S. Mozaffari, L. Balicas, F. F. Balakirev, V. B. Prakapenka, S. Chariton, D. A. Knyazev, E. Greenberg, and M. I. Eremets, Nat Commun 12, 5075 (2021).
- Troyan et al. (2021) I. A. Troyan, D. V. Semenok, A. G. Kvashnin, A. V. Sadakov, O. A. Sobolevskiy, V. M. Pudalov, A. G. Ivanova, V. B. Prakapenka, E. Greenberg, A. G. Gavriliuk, I. S. Lyubutin, V. V. Struzhkin, A. Bergara, I. Errea, R. Bianco, M. Calandra, F. Mauri, L. Monacelli, R. Akashi, and A. R. Oganov, Advanced Materials 33, 2006832 (2021).
- Li et al. (2022) Z. Li, X. He, C. Zhang, X. Wang, S. Zhang, Y. Jia, S. Feng, K. Lu, J. Zhao, J. Zhang, B. Min, Y. Long, R. Yu, L. Wang, M. Ye, Z. Zhang, V. Prakapenka, S. Chariton, P. A. Ginsberg, J. Bass, S. Yuan, H. Liu, and C. Jin, Nat Commun 13, 2863 (2022).
- Ma et al. (2022) L. Ma, K. Wang, Y. Xie, X. Yang, Y. Wang, M. Zhou, H. Liu, X. Yu, Y. Zhao, H. Wang, G. Liu, and Y. Ma, Phys. Rev. Lett. 128, 167001 (2022).
- Song et al. (2023) Y. Song, J. Bi, Y. Nakamoto, K. Shimizu, H. Liu, B. Zou, G. Liu, H. Wang, and Y. Ma, Phys. Rev. Lett. 130, 266001 (2023).
- Cross et al. (2024) S. Cross, J. Buhot, A. Brooks, W. Thomas, A. Kleppe, O. Lord, and S. Friedemann, Phys. Rev. B 109, L020503 (2024).
- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Z. Liu, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature 621, 493 (2023).
- Sakakibara et al. (2024) H. Sakakibara, M. Ochi, H. Nagata, Y. Ueki, H. Sakurai, R. Matsumoto, K. Terashima, K. Hirose, H. Ohta, M. Kato, Y. Takano, and K. Kuroki, Phys. Rev. B 109, 144511 (2024).
- Liu et al. (2022) X. Liu, P. Jiang, Y. Wang, M. Li, N. Li, Q. Zhang, Y. Wang, Y.-L. Li, and W. Yang, Phys. Rev. B 105, 224511 (2022).
- Zhang et al. (2022) C. Zhang, X. He, C. Liu, Z. Li, K. Lu, S. Zhang, S. Feng, X. Wang, Y. Peng, Y. Long, R. Yu, L. Wang, V. Prakapenka, S. Chariton, Q. Li, H. Liu, C. Chen, and C. Jin, Nat Commun 13, 5411 (2022).
- Ying et al. (2023) J. Ying, S. Liu, Q. Lu, X. Wen, Z. Gui, Y. Zhang, X. Wang, J. Sun, and X. Chen, Phys. Rev. Lett. 130, 256002 (2023).
- Buckel and Hilsch (1954) W. Buckel and R. Hilsch, Z. Physik 138, 109 (1954).
- Abeles et al. (1966) B. Abeles, R. W. Cohen, and G. W. Cullen, Phys. Rev. Lett. 17, 632 (1966).
- Garland et al. (1968) J. W. Garland, K. H. Bennemann, and F. M. Mueller, Phys. Rev. Lett. 21, 1315 (1968).
- Strongin et al. (1968) M. Strongin, O. F. Kammerer, J. E. Crow, R. D. Parks, D. H. Douglass, and M. A. Jensen, Phys. Rev. Lett. 21, 1320 (1968).
- Tian et al. (2005) M. Tian, J. Wang, J. S. Kurtz, Y. Liu, M. H. W. Chan, T. S. Mayer, and T. E. Mallouk, Phys. Rev. B 71, 104521 (2005).
- Beutel et al. (2016) M. H. Beutel, N. G. Ebensperger, M. Thiemann, G. Untereiner, V. Fritz, M. Javaheri, J. Nägele, R. Rösslhuber, M. Dressel, and M. Scheffler, Supercond. Sci. Technol. 29, 085011 (2016).
- Bang et al. (2019) W. Bang, T. D. Morrison, K. D. D. Rathnayaka, I. F. Lyuksyutov, D. G. Naugle, and W. Teizer, Thin Solid Films 676, 138 (2019).
- Lozano et al. (2019) D. P. Lozano, S. Couet, C. Petermann, G. Hamoir, J. K. Jochum, T. Picot, E. Menéndez, K. Houben, V. Joly, V. A. Antohe, M. Y. Hu, B. M. Leu, A. Alatas, A. H. Said, S. Roelants, B. Partoens, M. V. Milošević, F. M. Peeters, L. Piraux, J. Van de Vondel, A. Vantomme, K. Temst, and M. J. Van Bael, Phys. Rev. B 99, 064512 (2019).
- Knorr and Barth (1970) K. Knorr and N. Barth, Solid State Communications 8, 1085 (1970).
- Houben et al. (2017) K. Houben, S. Couet, M. Trekels, E. Menéndez, T. Peissker, J. W. Seo, M. Y. Hu, J. Y. Zhao, E. E. Alp, S. Roelants, B. Partoens, M. V. Milošević, F. M. Peeters, D. Bessas, S. A. Brown, A. Vantomme, K. Temst, and M. J. Van Bael, Phys. Rev. B 95, 155413 (2017).
- Houben et al. (2020) K. Houben, J. K. Jochum, S. Couet, E. Menéndez, T. Picot, M. Y. Hu, J. Y. Zhao, E. E. Alp, A. Vantomme, K. Temst, and M. J. Van Bael, Sci Rep 10, 5729 (2020).
- Barnett et al. (2004) J. D. Barnett, V. E. Bean, and H. T. Hall, Journal of Applied Physics 37, 875 (2004).
- Salamat et al. (2013) A. Salamat, R. Briggs, P. Bouvier, S. Petitgirard, A. Dewaele, M. E. Cutler, F. Corà, D. Daisenberger, G. Garbarino, and P. F. McMillan, Phys. Rev. B 88, 104104 (2013).
- Deffrennes et al. (2022) G. Deffrennes, P. Faure, F. Bottin, J.-M. Joubert, and B. Oudot, Journal of Alloys and Compounds 919, 165675 (2022).
- Wittig (1966) J. Wittig, Z. Physik 195, 228 (1966).
- Matsumoto et al. (2016) R. Matsumoto, Y. Sasama, M. Fujioka, T. Irifune, M. Tanaka, T. Yamaguchi, H. Takeya, and Y. Takano, Review of Scientific Instruments 87, 076103 (2016).
- Matsumoto et al. (2017) R. Matsumoto, T. Irifune, M. Tanaka, H. Takeya, and Y. Takano, Jpn. J. Appl. Phys. 56, 05FC01 (2017).
- Matsumoto et al. (2018) R. Matsumoto, A. Yamashita, H. Hara, T. Irifune, S. Adachi, H. Takeya, and Y. Takano, Appl. Phys. Express 11, 053101 (2018).
- Mao et al. (2008) H. K. Mao, P. M. Bell, J. W. Shaner, and D. J. Steinberg, Journal of Applied Physics 49, 3276 (2008).
- Akahama and Kawamura (2004) Y. Akahama and H. Kawamura, Journal of Applied Physics 96, 3748 (2004).
- Takano et al. (2004) Y. Takano, M. Nagao, I. Sakaguchi, M. Tachiki, T. Hatano, K. Kobayashi, H. Umezawa, and H. Kawarada, Applied Physics Letters 85, 2851 (2004).
- Adachi et al. (2020) S. Adachi, R. Matsumoto, S. Yamamoto, T. D. Yamamoto, K. Terashima, Y. Saito, M. Esparza Echevarria, P. Baptista de Castro, P. Song, S. Iwasaki, H. Takeya, and Y. Takano, Applied Physics Letters 116, 223506 (2020).
- Matsumoto et al. (2021) R. Matsumoto, S. Nakano, S. Yamamoto, and Y. Takano, Jpn. J. Appl. Phys. 60, 090902 (2021).
- Mito et al. (2001) M. Mito, M. Hitaka, T. Kawae, K. Takeda, T. Kitai, and N. Toyoshima, Jpn. J. Appl. Phys. 40, 6641 (2001).
- Mito et al. (2014) M. Mito, T. Imakyurei, H. Deguchi, K. Matsumoto, H. Hara, T. Ozaki, H. Takeya, and Y. Takano, J. Phys. Soc. Jpn. 83, 023705 (2014).
- Mito et al. (2016) M. Mito, H. Goto, H. Matsui, H. Deguchi, K. Matsumoto, H. Hara, T. Ozaki, H. Takeya, and Y. Takano, J. Phys. Soc. Jpn. 85, 024711 (2016).
- Mito et al. (2017) M. Mito, K. Ogata, H. Goto, K. Tsuruta, K. Nakamura, H. Deguchi, T. Horide, K. Matsumoto, T. Tajiri, H. Hara, T. Ozaki, H. Takeya, and Y. Takano, Phys. Rev. B 95, 064503 (2017).
- Abdel-Hafiez et al. (2018) M. Abdel-Hafiez, Y. Zhao, Z. Huang, C.-w. Cho, C. H. Wong, A. Hassen, M. Ohkuma, Y.-W. Fang, B.-J. Pan, Z.-A. Ren, A. Sadakov, A. Usoltsev, V. Pudalov, M. Mito, R. Lortz, C. Krellner, and W. Yang, Phys. Rev. B 97, 134508 (2018).
- Mito et al. (2019) M. Mito, Y. Kitamura, T. Tajiri, K. Nakamura, R. Shiraishi, K. Ogata, H. Deguchi, T. Yamaguchi, N. Takeshita, T. Nishizaki, K. Edalati, and Z. Horita, Journal of Applied Physics 125, 125901 (2019).
- Irifune et al. (2003) T. Irifune, A. Kurio, S. Sakamoto, T. Inoue, and H. Sumiya, Nature 421, 599 (2003).
- Dolan and Silcox (1973) G. J. Dolan and J. Silcox, Phys. Rev. Lett. 30, 603 (1973).
- Ohkuma et al. (2023) M. Ohkuma, R. Matsumoto, and Y. Takano, Appl. Phys. Express 16, 043004 (2023).
- Helfand and Werthamer (1966) E. Helfand and N. R. Werthamer, Phys. Rev. 147, 288 (1966).
- Werthamer et al. (1966) N. R. Werthamer, E. Helfand, and P. C. Hohenberg, Phys. Rev. 147, 295 (1966).
- Baumgartner et al. (2013) T. Baumgartner, M. Eisterer, H. W. Weber, R. Flükiger, C. Scheuerlein, and L. Bottura, Supercond. Sci. Technol. 27, 015005 (2013).
- Zhang et al. (2016) Y. Zhang, C. H. Wong, J. Shen, S. T. Sze, B. Zhang, H. Zhang, Y. Dong, H. Xu, Z. Yan, Y. Li, X. Hu, and R. Lortz, Sci Rep 6, 32963 (2016).
- Shao et al. (2018) Z. Shao, Z. Zhang, H. Yuan, H. Sun, Y. Cao, X. Zhang, S. Li, H. Gedeon, T. Xiang, Q.-K. Xue, and M. Pan, Science Bulletin 63, 1332 (2018).
- Glover (1967) R. E. Glover, Physics Letters A 25, 542 (1967).
- Aslamasov and Larkin (1968) L. G. Aslamasov and A. I. Larkin, Physics Letters A 26, 238 (1968).
- Skocpol and Tinkham (1975) W. J. Skocpol and M. Tinkham, Rep. Prog. Phys. 38, 1049 (1975).
- Sóñora et al. (2019) D. Sóñora, C. Carballeira, J. J. Ponte, F. Vidal, T. Grenet, and J. Mosqueira, Phys. Rev. B 100, 104509 (2019).
- Hsieh et al. (2019) S. Hsieh, P. Bhattacharyya, C. Zu, T. Mittiga, T. J. Smart, F. Machado, B. Kobrin, T. O. Höhn, N. Z. Rui, M. Kamrani, S. Chatterjee, S. Choi, M. Zaletel, V. V. Struzhkin, J. E. Moore, V. I. Levitas, R. Jeanloz, and N. Y. Yao, Science 366, 1349 (2019).
- Lesik et al. (2019) M. Lesik, T. Plisson, L. Toraille, J. Renaud, F. Occelli, M. Schmidt, O. Salord, A. Delobbe, T. Debuisschert, L. Rondin, P. Loubeyre, and J.-F. Roch, Science 366, 1359 (2019).
- Umeo (2016) K. Umeo, Review of Scientific Instruments 87, 063901 (2016).
- Cao et al. (2023) Z.-Y. Cao, H. Jang, S. Choi, J. Kim, S. Kim, J.-B. Zhang, A. S. Sharbirin, J. Kim, and T. Park, NPG Asia Mater 15, 1 (2023).

### v2

@@ Abstract
We investigated the pressure effects of a superconductivity on thin films of Sn. Elemental superconductor Sn with a body-centered tetragonal structure, -Sn, exhibits superconductivity below the superconducting transition temperature ( K) at ambient pressure. of Sn increases with lowering dimension such as in thin film and nanowire growth, or by high-pressure application. For thin films, exhibits a slight increase up to approximately 4 K compared to the bulk value, attributable to the crystalline size and lattice disorder. By applying pressure on a bulk Sn, initially decreases from 3.72 K as the pressure increases. Further increasing pressure up to 10 GPa, increases to 5.3 K with the structural transformation. However, the combination of these effects on thin films of Sn, namely, thin-film growth and pressure effects, remains underexplored. In this study, we combined film-growth and pressure-application techniques to further increase using a diamond anvil cell with boron-doped diamond electrodes. The drop of the electrical resistance suggesting the onset of on the thin film reached above 6 K in -Sn phase. Further, the upper critical magnetic field was drastically enhanced. Atomic force microscopy suggests that the refinement of the grain size of the thin film under the non-hydrostatic pressure conditions contributes to stabilizing the higher of -Sn.
@@ III.3 Non-hydrostatic pressure on thin film
Figures 4(a) and (b) show the dependence of for the thin film of Sn under various at 2.5 and 10.5 GPa under non-hydrostatic pressure condition. was observed at 1 T, whereas the upper critical magnetic field () was approximately 0.1 T under ambient pressure. Drastic enhancement was also observed in the phase at 10.5 GPa. As shown in Fig. 4(c), estimated using Werthamer–Helfand–Hohenberg (WHH) model reached several teslas at high pressures Helfand and Werthamer (1966); Werthamer et al. (1966); Baumgartner et al. (2014). Notably, the enhancement of -Sn under the non-hydrostatic pressure condition was observed reproducibly. Figure 5(b) shows the dependence of for the thin film of Sn with the other setup at 5 GPa. The optical image of the thin film of Sn is shown in Fig. 5(a). The decrease suggesting the superconducting transition was observed even at T.
There are two possible scenarios for the enhancement. The first is the enhancement of the flux-pinning force. The crystallographic defects introduced by non-hydrostatic pressure could serve as pinning centers, potentially leading to an increase in . The other is the shortening of the mean free path of the electrons and the resulting shortening of the coherence length. Considering that the thickness of the thin film was approximately 100 nm, the electrical resistivity of the normal state near was estimated to be approximately for ambient pressure measurement. This value is consistent with those reported in previous studies Bang et al. (2019); Ohkuma et al. (2023). In contrast, the electrical resistivities of the thin films for 2.5 GPa of Run 1 and 5 GPa of Run 2 were approximately and , respectively. The increase in the electrical resistivity near suggests that the mean free path of the electrons is shortened by application of pressure.
@@ III.4 Quasi-hydrostatic pressure on thin film
We performed high-pressure generation with better hydrostatic pressure condition using a liquid pressure transmitted medium, glycerol. Figure 6(a) shows the – of the thin film at 9 GPa under quasi-hydrostatic pressure conditions. The optical image of the thin film inside the DAC is shown in the inset of Fig. 6(a). The RRR was with in 2–3 in quasi-hydrostatic pressure measurements, which was slightly higher than those for the non-hydrostatic pressure condition. The drop suggesting superconducting transition was observed around 5 K. Figure 6(b) shows – under quasi-hydrostatic pressure condition between 2 and 8 K with warming process. The value was normalized using the value at 8 K. At 9 GPa, slightly decreased with decreasing around 5.3 K, suggesting the superconducting transition on -Sn. The drop was also observed around 3 K. The -Sn phase became significant on further pressure application. The onset of slightly increased at 12 GPa. We also decreased the pressure from 12 to 8 GPa. The decrease suggesting the superconducting transition of -Sn was slightly observed, whereas the -Sn was clearly observed at 8.5 GPa under the non-hydrostatic pressure condition. With further pressure decrease, -Sn vanished at 2 GPa.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant No. 19H02177, 20H05644 and JST-Mirai Program Grant No. JPMJMI17A2. This work was also supported by the Joint Usage/Research Center PRIUS, Ehime University, JapanData Availability Statement
The data that support the findings of this article are openly available Sasaki et al. (2025)